Hello I am still having trouble with this find the rate problem. If you could help me I have tried to find the time but it just does't work out to the right answer and I have the have the answer. But I just don't know how. I would be very grateful.
Jeremy walked 10 miled in one-half hour longer than it took Hugh to walk 7 miles. Jeremy walked at an average rate of one-half mile per hour faster than Hugh. Find the rates of both Jeremy and Hugh.
Thank You so much,
Tiffany

Answer by rwm(914)

So you want me to show you all of my work with the answer but you don't even share any of your work and the answer?
Jeremy walked faster and for more time
r=speed t= time d=distance
r*t=d
r and t are the speed and time for hugh
r+.5 and t+.5 are the speed and time for jeremy

(r+.5)*(t+.5)=10
r*t=7
solve for
r and r+.5
t and t+.5